OpenStudy (anonymous):

3. Solve: 0 = 6x2 − 10x − 4. (1 point) x = −1 over 3, x = −2 x = −1 over 3, x = 2 x = 1 over 3, x = −2 x = 1 over 3, x = 2

OpenStudy (anonymous):

\[6x^2-10x-4=0\]divide both sides by 2\[3x^2-5x-2=0\]use quadratic formula

OpenStudy (anonymous):

He's right about the equation, but don't use quadratic formula, because it is factorable normally. \[3x ^{2} - 5x - 2 = 0\] Can be written as: \[(3x+1)(x-2)\]

OpenStudy (anonymous):

So for your anwer, I would choose: x = -1 over 3 and, x = 2 the second option!

OpenStudy (anonymous):

@nj1202 is wright